# Elliptic boundary-value problems in some distribution spaces of generalized smoothness

### v1

@@ 2. Spaces of generalized smoothness
endowed with the norm $\|\cdot,l^{s,\varphi}_{p}(E)\|$, where $E:=L_{p}(\mathbb{R}^{d})$. According to [9, Theorem 2.5(ii)], every space $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ is a retract of $l^{s,\varphi}_{p}(L_{p}(\mathbb{R}^{d}))$, with the corresponding retraction operator $\mathcal{R}$ and coretraction operator $\mathcal{T}$ being independent (as mappings) of $s$, $\varphi$, and $p$. This means that we have certain bounded linear operators
@@
such that $\mathcal{R}\mathcal{T}$ is the identical operator. Hence,
$[B^{s_{0},\varphi_{0}}_{p_{0}}(\mathbb{R}^{d}),B^{s_{1},\varphi_{1}}_{p_{1}}(\mathbb{R}^{d})]_{\theta}$ (2.23)
@@
$\|w,l_{p}\{E_{k}\}\|^{p}:=\Biggl{(}\,\sum_{k=0}^{\infty}\|w_{k},E_{k}\|^{p}\Biggr{)}^{1/p}=\|w,l^{s,\varphi}_{p}(E)\|<\infty,$
@@
These equalities hold true with equalities of norms. The second equality is due to [62, Theorems 1.18.1 and Remark 1.18.4/1], the third one is valid because the complex interpolation functor is precise of type $\theta$ [62, Theorem 1.9.3(a)], and the forth one is due to [62, Theorem 1.18.4 and Remark 1.18.4/2].
@@
Thus, the spaces $B^{s,\varphi}_{p}(\Gamma)$ and $[B^{s_{0},\varphi_{0}}_{p_{0}}(\Gamma),B^{s_{1},\varphi_{1}}_{p_{1}}(\Gamma)]_{\theta}$ are retracts of the same space $(B^{s,\varphi}_{p}(\mathbb{R}^{d}))^{\varkappa}$ with the same retraction and coretraction operators. Therefore, equality (2.21) holds true up to equivalence of norms in the $W=\Gamma$ case.
@@
Let $1<p<\infty$, $s>1/p$, and $\varphi\in\Upsilon$. Then the trace mapping $R_{\Gamma}:u\mapsto u\!\upharpoonright\!\Gamma$, where $u\in C^{\infty}(\overline{\Omega})$, extend uniquely (by continuity) to a bounded linear operator
@@
As usual, $u\!\upharpoonright\!\Gamma$ denotes the restriction of the function $u$ to $\Gamma$.
@@
This result is known for $\varphi(\cdot)\equiv 1$ (e.g., it contains in [62, Lemma 5.4.4]) and for $p=2$ [44, Theorem 3.5]. We will deduce it from this fact by the interpolation Theorems 2.3 and 2.7. We need to consider the cases $21<p<$ 2 and $2<p<\infty$ separately, which is caused by the restriction $s>1/p$.
We first consider the case where $21<p<$ 2. Choose real numbers $p_{1}>1$ and $s_{1}$ so that $1/p<1/p_{1}<s_{1}<s$, and note that $21<p_{1}<p<$ 2. We then define the parameters $\theta$, $s_{0}$, and $\varphi_{0}$ according to Theorem 2.3 and see that $s_{0}>s>1/p>1/2$ in the case under consideration. Therefore, we have the following two pairs of bounded linear operators:
@@
[…]
@@
where $s>1/p$. This isomorphism is induced by the Dirichlet problem for the elliptic equation $u-\Delta u=f$ in the half-space +$\mathbb{R}^{n}_{+}$ (as usual, $\Delta$ is the Laplace operator), we naturally identifying $\mathbb{R}^{n}_{0}$ with $\mathbb{R}^{n-1}$. The operator $S_{0}$ can be defined as the restriction of the inverse of (2.31) to the space $\{0\}\times B^{s-1/p}_{p}(\mathbb{R}^{n-1})$. Thus, $B^{s-1/p,\varphi}_{p}(\mathbb{R}^{n-1})$ is the trace space for $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ and $H^{s,\varphi}_{p}(\mathbb{R}^{n}_{+})$ whenever $1<p<\infty$, $s>1/p$, and $\varphi\in\Upsilon$.
@@
A necessary and sufficient condition for the Sobolev space of generalized smoothness over $\mathbb{R}^{n}$ under which every function from the space has a trace on a hyperplane was obtained by Hörmander [23, Theorem 2.2.8] in the case where the integral-exponent $p$ equals $2$ (see also Volevich and Paneah’s paper [64, Theorems 6.1 and 6.2]) and by Goldman [17, Theorem 1] for each $p\in(1,\infty)$. The result by Hörmander concerns some (generally, anisotropic) spaces of generalized smoothness with the integral-exponent $p\in[1,\infty]$ and also describes their trace spaces. These spaces coincide with Sobolev spaces of generalized smoothness if and only if $p=2$. Goldman [17, Theorem 2] indicated a Besov space of generalized smoothness that contains these traces but did not assert that it is the narrowest space. The trace spaces were found by Kalyabin [27, § 1] for certain anisotropic Sobolev spaces of generalized smoothness; however, the class of these spaces does not contain isotropic ones. For isotropic Sobolev spaces of generalized smoothness, the description of the trace spaces contains in Kalyabin’s later result [28, Section 1, Main Theorem] devoted to Triebel–Lizorkin spaces of generalized smoothness (both isotropic and anisotropic). Goldman [17, Corollary 3] and Merucci [39, Theorem 15] indicated classes of Sobolev spaces of generalized smoothness for which the traces belong to the Besov space whose smoothness index equals $t^{-1/p}\omega(t)$ provided that the function $\omega$ is the smoothness index of the Sobolev space. However, it was not showed that this Besov space is the trace space (see, e.g., [39, Remark 5]). The spaces $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ for which $s>1/p$ and $\varphi\in\Upsilon$ belong to these classes, with $\omega(t)\equiv t^{s}\varphi(t)$.
@@
$\int_{1}^{\infty}\frac{dt}{t\,\varphi^{p^{\prime}}(t)}<\infty,$ (2.32)
which follows from [17, Theorem 1]; as usual, $11/p+1/p^{\prime}=$ 1. If (2.32) holds true, then the trace space for $H^{1/p,\varphi}_{p}(\Omega)$ differs from $B^{0,\varphi}_{p}(\Gamma)$. This follows, e.g., from [23, Theorem 2.2.8] or [64, Theorem 6.2] in the $p=2$ case, as is shown in [41, Theorem 3.8 b)]. The same concerns the trace mapping $R_{0}$ indicated in Remark 2.11. Running ahead, note that the Fredholm property of elliptic boundary-value problems in $\Omega$ will not hold true if we suppose (2.32) and take $H^{1/p,\varphi}_{p}(\Omega)$ as a solution space of these problems. This is explained in [43, Section 9.3] in the $p=2$ case.
@@ 3. Applications to elliptic problems
Recall that the ellipticity of this problem means that the PDO $A$ is properly elliptic on $\overline{\Omega}$ and that the collection $B$ of boundary PDOs satisfy the Lopatinskii condition with respect to $A$ on $\Gamma$ (see, e.g., [3, Section 1.2]). (If $n\geq 3$, then the elipticity of $A$ implies its proper ellipticity.)
@@
[…]
@@
Proof of the Theorem 3.1.
If $\varphi(\cdot)\equiv 1$, then this theorem is a classical result (see, e.g., [62, Theorem 5.5.2] in the case where the elliptic problem is regular and $s\geq 2q$; this result is also contained in [26, Theorem 5.2]). If $p=2$ and $\varphi\in\Upsilon$ (the case of Hilbert spaces), Theorem 3.1 is contained in [41, Theorem 4.1] (if $m\leq 2q-1$) and [32, Theorem 1] (if $m\geq 2q$). Let $N$ and $\varkappa$ respectively denote the kernel and index of the Fredholm operator (3.4) provided that $1<p<\infty$ and $\varphi(\cdot)\equiv 1$ or that $p=2$ and $\varphi\in\Upsilon$, with $N$ and $\varkappa$ not depending on $\varphi$ and $p$ indicated. Let $M$ satisfy the conclusion of Theorem 3.1 for these $\varphi$ and $p$. We need to treat the cases where $21<p<$ 2 and where $p>2$ separately.
We first consider the $21<p<$ 2 case. Choose numbers $p_{1}$ and $s_{1}$ so that $1<p_{1}<p$ and $m+1/p<m+1/p_{1}<s_{1}<s$. Define parameters $\theta\in(0,1)$, $s_{0}$, and $\varphi_{0}\in\Upsilon$ according to the hypotheses of Theorem 2.3. Note that $s_{0}>s_{1}>m+1/p_{1}>m+1/2$ and $2>p>p_{1}$. Hence, we have the dense continuous embeddings
@@
Moreover, we have the Fredholm bounded operators
@@
$[H^{s_{0},\varphi_{0}}_{2}(\Omega),H^{s_{1}}_{p_{1}}(\Omega)]_{\theta}=H^{s,\varphi}_{p}(\Omega)$ (3.9)
@@
because $M\subset\Xi^{s-2q,\varphi}_{p}(\Omega,\Gamma)$; i.e. (3.5) holds true. Since $C^{\infty}(\overline{\Omega})$ is dense in $H^{s,\varphi}_{p}(\Omega)$, this operator is the extension by continuity of mapping (3.3). We have proved Theorem 3.1 in the case where $21<p<$ 2.
@@
We also have the Fredholm bounded operators
@@
$[H^{s_{0}}_{p}(\Omega),H^{s_{1},\varphi_{1}}_{p}(\Omega)]_{\theta}=H^{s,\varphi}_{p}(\Omega)$
@@
$[\Xi^{s_{0}-2q}_{p}(\Omega,\Gamma),\Xi^{s_{1}-2q,\varphi_{1}}_{p}(\Omega,\Gamma)]_{\theta}=\Xi^{s-2q,\varphi}_{p}(\Omega,\Gamma).$
@@
Let us study local properties (up to the boundary $\Gamma$) of the generalized solutions to the elliptic problem. To this end, we introduce local versions of distribution spaces appeared in (3.4). Let $U$ be an open set in $\mathbb{R}^{n}$ such that $\Omega_{0}:=\Omega\cap U\neq\emptyset$, and put $\Gamma_{0}:=\Gamma\cap U$ (the $\Gamma_{0}=\emptyset$ case is admissible). Given $\sigma\in\mathbb{R}$, $\varphi\in\Upsilon$, and $p\in(1,\infty)$, we let $H^{\sigma,\varphi}_{p,\mathrm{loc}}(\Omega_{0},\Gamma_{0})$ denote the linear space of all distributions $u\in\mathcal{D}^{\prime}(\Omega)$ such that $\chi u\in H^{\sigma,\varphi}_{p}(\Omega)$ whenever $\chi\in C^{\infty}(\overline{\Omega})$ and $\mathrm{supp}\,\chi\subset\Omega_{0}\cup\Gamma_{0}$. Analogously, $B^{\sigma,\varphi}_{p,\mathrm{loc}}(\Gamma_{0})$ denotes the linear space of all distributions $v\in\mathcal{D}^{\prime}(\Gamma)$ such that $\chi v\in B^{\sigma,\varphi}_{p}(\Gamma)$ whenever $\chi\in C^{\infty}(\Gamma)$ and $\mathrm{supp}\,\chi\subset\Gamma_{0}$. The operator of the multiplication by any function $\chi\in C^{\infty}(\overline{\Omega})$ (resp., $\chi\in C^{\infty}(\Gamma)$) is bounded on $H^{\sigma,\varphi}_{p}(\Omega)$ (resp., $B^{\sigma,\varphi}_{p}(\Gamma)$), which is known in the $\varphi(\cdot)\equiv 1$ case and then extends by interpolation (Theorems 2.3, 2.7 and Propositions 2.1, 2.8) to the $\varphi\in\Upsilon$ case. Hence, $H^{\sigma,\varphi}_{p}(\Omega)\subset H^{\sigma,\varphi}_{p,\mathrm{loc}}(\Omega_{0},\Gamma_{0})$ and $B^{\sigma,\varphi}_{p}(\Gamma)\subset B^{\sigma,\varphi}_{p,\mathrm{loc}}(\Gamma_{0})$ If $\Omega_{0}=\Omega$ and $\Gamma_{0}=\Gamma$, then these inclusions become equalities.
@@
due to (3.5); i.e., $(f,g)=(A,B)v$ for certain $v\in H^{s,\varphi}_{p}(\Omega)$. Hence, the distribution $u-v\in H^{s-\varepsilon}_{p}(\Omega)$ satisfies $(A,B)(u-v)=0$, which yields $w:=u-v\in C^{\infty}(\overline{\Omega})$ by Theorem 3.1. Thus, $u=v+w\in H^{s,\varphi}_{p}(\Omega)$. We have proved Theorem 3.3 in the case where $\Omega_{0}=\Omega$ and $\Gamma_{0}=\Gamma$, which will help to treat the general case.
@@
Suppose that premise of the application is true. We consider an arbitrary function $\chi\in\Psi$ and choose a function $\eta\in\Psi$ so that $\eta=1$ in the neighbourhood of $\mathrm{supp}\,\chi$. By hypothesis (3.13), we have the inclusion $\chi F\in\Xi^{s-2q,\varphi}_{p}(\Omega,\Gamma)$ where $F:=(A,B)u$. Permuting the differential operator $(A,B)$ an the operator of the multiplication by $\chi$, we obtain
@@
Here, $(A^{\prime},B^{\prime})$ is a differential operator of the form $(A,B)$ with infinitely smooth coefficients such that the orders of all its components are less (at least by $1$) than the orders of the corresponding components of the operator $(A,B)$. (Of course, if $\mathrm{ord}\,B_{j}=0$ [i.e., $B_{j}$ is an operator of the multiplication by a function of class $C^{\infty}(\Gamma)$], then the corresponding operator $B_{j}^{\prime}=0$ and we may assume that $\mathrm{ord}\,B_{j}^{\prime}=-1$.)
@@
whatever $p>1$, $s>m-1+1/p$, and $\varphi\in\Upsilon$. This is deduced from the $\varphi(\cdot)\equiv 1$ case by interpolation in an analogous way to that used for (2.26) in the proof of Theorem 2.10, with the condition $s>m-1+1/p$ being stipulated by the fact that the orders of all components of $B^{\prime}$ are less than or equal to $m-1$.
@@
We supplement the last theorem with the corresponding a priory estimate of the solution $u$.
@@
If $\Omega_{0}=\Omega$ and $\Gamma_{0}=\Gamma$, then we may take $\chi(t)\equiv\eta(t)\equiv 1$ in (3.20), which means the global a priory estimate of $u$. In this case, estimate (3.20) follows by Peetre’s lemma [53, Lemma 3] from Theorem 3.1 (from the fact that operator (3.4) has finite-dimensional kernel and closed range) and the compact embedding $H^{s,\varphi}_{p}(\Omega)\hookrightarrow H^{s-1,\varphi}_{p}(\Omega)$ (in fact, the continuity of this embedding is enough). In the general case, this estimate is deduced from its global version with the help of (3.15) where $F=(A,B)u$. Namely:
@@
for a certain function parameter $\varphi\in\Upsilon$ subject to
$\int\limits_{1}^{\infty}\frac{dt}{t\,\varphi^{r}(t)}<\infty,\quad\mbox{with}\quad\frac{1}{p}+\frac{1}{r}=1.$ (3.22)
Then $u\in C^{\ell}(\Omega_{0}\cup\Gamma_{0})$.

### v2

@@ 2. Spaces of generalized smoothness
endowed with the norm $\|\cdot,l^{s,\varphi}_{p}(E)\|$, where $E:=L_{p}(\mathbb{R}^{d})$. According to [9, Theorem 2.5(ii)], every space $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ is a retract of $l^{s,\varphi}_{p}(L_{p}(\mathbb{R}^{d}))$, where the retraction operator $\mathcal{R}$ and the coretraction operator $\mathcal{T}$ can be chosen independent of $s$, $\varphi$, and $p$. This means that we have certain bounded linear operators
@@
such that $\mathcal{R}\mathcal{T}$ is the identity mapping. Hence,
$B^{s_{0},\varphi_{0}}_{p_{0}}(\mathbb{R}^{d}),B^{s_{1},\varphi_{1}}_{p_{1}}(\mathbb{R}^{d})]_{\theta}$ (2.23)
@@
$\|w,l_{p}\{E_{k}\}\|:=\Biggl(\,\sum_{k=0}^{\infty}\|w_{k},E_{k}\|^{p}\Biggr)^{1/p}=\|w,l^{s,\varphi}_{p}(E)\|<\infty,$
@@
These equalities hold true with equalities of norms. The second equality is due to [62, Theorems 1.18.1 and Remark 1.18.4/1], the third one is valid because the complex interpolation functor is exact of type $\theta$ [62, Theorem 1.9.3(a)], and the fourth one is due to [62, Theorem 1.18.4 and Remark 1.18.4/2].
@@
Thus, the spaces $B^{s,\varphi}_{p}(\Gamma)$ and $B^{s_{0},\varphi_{0}}_{p_{0}}(\Gamma),B^{s_{1},\varphi_{1}}_{p_{1}}(\Gamma)]_{\theta}$ are retracts of the same space $(B^{s,\varphi}_{p}(\mathbb{R}^{d}))^{\varkappa}$ with the same retraction and coretraction operators. Therefore, equality (2.21) holds true up to equivalence of norms in the $W=\Gamma$ case.
@@
Let $1<p<\infty$, $s>1/p$, and $\varphi\in\Upsilon$. Then the trace mapping $R_{\Gamma}:u\mapsto u\!\upharpoonright\!\Gamma$, where $u\in C^{\infty}(\overline{\Omega})$, extends uniquely (by continuity) to a bounded linear operator
@@
As usual, $u\!\upharpoonright\!\Gamma$ denotes the restriction of the function $u$ to $\Gamma$, and $C^{\infty}(\overline{\Omega})$ stands for the set of the restrictions of all functions $w\in C^{\infty}(\mathbb{R}^{n})$ to $\overline{\Omega}$. Since $\Omega$ is bounded, $C^{\infty}(\overline{\Omega})=C^{\infty}_{0}(\overline{\Omega})$ is a dense subset of $H^{s,\varphi}_{p}(\Omega)$.
@@
This result is known for $\varphi(\cdot)\equiv 1$ (e.g., it is contained in [62, Lemma 5.4.4]) and for $p=2$ [44, Theorem 3.5]. We will deduce it from this fact by the interpolation Theorems 2.3 and 2.7. We need to consider the cases $1<p<2$ and $2<p<\infty$ separately, which is caused by the restriction $s>1/p$.
We first consider the case where $1<p<2$. Choose real numbers $p_{1}>1$ and $s_{1}$ so that $1/p<1/p_{1}<s_{1}<s$, and note that $1<p_{1}<p<2$. We then define the parameters $\theta$, $s_{0}$, and $\varphi_{0}$ according to Theorem 2.3 and see that $s_{0}>s>1/p>1/2$ in the case under consideration. Therefore, we have the following two pairs of bounded linear operators:
@@
hold true up to equivalence of norms. Applying the complex interpolation with the parameter $\theta$ to these pairs of operators, we obtain bounded operators (2.26) and (2.27) such that (2.26) is an extension by continuity of the mapping $C^{\infty}(\overline{\Omega})\ni u\mapsto u\!\upharpoonright\!\Gamma$ and that (2.27) is a right-inverse of (2.26). Hence, the operator (2.26) is onto, and (2.28) holds true.
@@
where $s>1/p$. This isomorphism is induced by the Dirichlet problem for the elliptic equation $u-\Delta u=f$ in the half-space $\mathbb{R}^{n}_{+}$ (as usual, $\Delta$ is the Laplace operator), we naturally identifying $\mathbb{R}^{n}_{0}$ with $\mathbb{R}^{n-1}$. The operator $S_{0}$ can be defined as the restriction of the inverse of (2.31) to the space $\{0\}\times B^{s-1/p}_{p}(\mathbb{R}^{n-1})$. Thus, $B^{s-1/p,\varphi}_{p}(\mathbb{R}^{n-1})$ is the trace space for $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ and $H^{s,\varphi}_{p}(\mathbb{R}^{n}_{+})$ whenever $1<p<\infty$, $s>1/p$, and $\varphi\in\Upsilon$.
@@
A necessary and sufficient condition for the Sobolev space of generalized smoothness over $\mathbb{R}^{n}$ under which every function from the space has a trace on a hyperplane was obtained by Hörmander [23, Theorem 2.2.8] in the case where the integrability exponent $p$ equals $2$ (see also Volevich and Paneah’s paper [64, Theorems 6.1 and 6.2]) and by Goldman [17, Theorem 1] for each $p\in(1,\infty)$. The result by Hörmander concerns some (generally, anisotropic) spaces of generalized smoothness with the integrability exponent $p\in[1,\infty]$ and also describes their trace spaces. These spaces coincide with Sobolev spaces of generalized smoothness if and only if $p=2$. Goldman [17, Theorem 2] indicated a Besov space of generalized smoothness that contains these traces but did not assert that it is the narrowest space. The trace spaces were found by Kalyabin [27, § 1] for certain anisotropic Sobolev spaces of generalized smoothness; however, the class of these spaces does not contain isotropic ones. For isotropic Sobolev spaces of generalized smoothness, the description of the trace spaces contains in Kalyabin’s later result [28, Section 1, Main Theorem] devoted to Triebel–Lizorkin spaces of generalized smoothness (both isotropic and anisotropic). Goldman [17, Corollary 3] and Merucci [39, Theorem 15] indicated classes of Sobolev spaces of generalized smoothness for which the traces belong to the Besov space whose smoothness index equals $t^{-1/p}\omega(t)$ provided that the function $\omega$ is the smoothness index of the Sobolev space. However, it was not shown that this Besov space is the trace space (see, e.g., [39, Remark 5]). The spaces $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ for which $s>1/p$ and $\varphi\in\Upsilon$ belong to these classes, with $\omega(t)\equiv t^{s}\varphi(t)$.
@@
$\int\limits_{1}^{\infty}\frac{dt}{t\,\varphi^{p^{\prime}}(t)}<\infty,$ (2.32)
which follows from [17, Theorem 1]; as usual, $1/p+1/p^{\prime}=1$. If (2.32) holds true, then the trace space for $H^{1/p,\varphi}_{p}(\Omega)$ differs from $B^{0,\varphi}_{p}(\Gamma)$. This follows, e.g., from [23, Theorem 2.2.8] or [64, Theorem 6.2] in the $p=2$ case, as is shown in [41, Theorem 3.8 b)]. The same concerns the trace mapping $R_{0}$ indicated in Remark 2.11. Running ahead, note that the Fredholm property of elliptic boundary-value problems in $\Omega$ will not hold true if we suppose (2.32) and take $H^{1/p,\varphi}_{p}(\Omega)$ as a solution space of these problems. This is explained in [43, Section 9.3] in the $p=2$ case.
@@ 3. Applications to elliptic problems
Recall that the ellipticity of this problem means that the PDO $A$ is properly elliptic on $\overline{\Omega}$ and that the collection $B$ of boundary PDOs satisfies the Lopatinskii condition with respect to $A$ on $\Gamma$ (see, e.g., [3, Section 1.2]). (If $n\geq 3$, then the elipticity of $A$ implies its proper ellipticity.)
@@
Let X=[X $X_{\,0},X_{1}$] and Y=[Y $Y_{\,0},Y_{1}$] be ordered pairs of Banach spaces such that the dense continuous embeddings $X_{1}\hookrightarrow X_{0}$ and $Y_{1}\hookrightarrow Y_{0}$ (or $X_{0}\hookrightarrow X_{1}$ and $Y_{0}\hookrightarrow Y_{1}$) hold true. Suppose that a linear mapping $T$ is given on $X_{\,0}$ (resp. $X_{\,1}$) and defines bounded Fredholm operators $T:X_{0}\rightarrow Y_{0}$ and $T:X_{1}\rightarrow Y_{1}$ with common kernel $N$ and the same index $\varkappa$. Then for any interpolation parameter $\theta\in(0,1)$ the bounded operator $T:[X_{\,0},X_{1}]_{\theta}\rightarrow[Y_{\,0},Y_{1}]_{\theta}$ is a Fredholm operator with the kernel $N$, range $Y_{\,0},Y_{1}]_{\theta}\cap T(X_{\,0})$ (resp. $Y_{\,0},Y_{1}]_{\theta}\cap T(X_{\,1})$) and the same index $\varkappa$.
@@
Proof of Theorem 3.1.
If $\varphi(\cdot)\equiv 1$, then this theorem is a classical result (see, e.g., [62, Theorem 5.5.2] in the case where the collection $(A,B_{1},\ldots,B_{q})$ is regularly elliptic [62, Definition 5.2.1/4] and $s\geq 2q$; this result is also contained in [26, Theorem 5.2]). If $p=2$ and $\varphi\in\Upsilon$ (the case of Hilbert spaces), Theorem 3.1 is contained in [41, Theorem 4.1] (if $m\leq 2q-1$) and [32, Theorem 1] (if $m\geq 2q$). Let $N$ and $\varkappa$ respectively denote the kernel and index of the Fredholm operator (3.4) provided that $1<p<\infty$ and $\varphi(\cdot)\equiv 1$ or that $p=2$ and $\varphi\in\Upsilon$, with $N$ and $\varkappa$ not depending on $\varphi$ and $p$ indicated. Let $M$ satisfy the conclusion of Theorem 3.1 for these $\varphi$ and $p$. We need to treat the cases where $1<p<2$ and where $p>2$ separately.
We first consider the $1<p<2$ case. Choose numbers $p_{1}$ and $s_{1}$ so that $1<p_{1}<p$ and $m+1/p<m+1/p_{1}<s_{1}<s$. Define parameters $\theta\in(0,1)$, $s_{0}$, and $\varphi_{0}\in\Upsilon$ according to the hypotheses of Theorem 2.3. Note that $s_{0}>s_{1}>m+1/p_{1}>m+1/2$ and $2>p>p_{1}$. Hence, we have the dense continuous embeddings
@@
Moreover, we have the bounded Fredholm operators
@@
$H^{s_{0},\varphi_{0}}_{2}(\Omega),H^{s_{1}}_{p_{1}}(\Omega)]_{\theta}=H^{s,\varphi}_{p}(\Omega)$ (3.9)
@@
because $M\subset\Xi^{s-2q,\varphi}_{p}(\Omega,\Gamma)$; i.e. (3.5) holds true. Since $C^{\infty}(\overline{\Omega})$ is dense in $H^{s,\varphi}_{p}(\Omega)$, this operator is the extension by continuity of mapping (3.3). We have proved Theorem 3.1 in the case where $1<p<2$.
@@
We also have the bounded Fredholm operators
@@
$H^{s_{0}}_{p}(\Omega),H^{s_{1},\varphi_{1}}_{p}(\Omega)]_{\theta}=H^{s,\varphi}_{p}(\Omega)$
@@
$\Xi^{s_{0}-2q}_{p}(\Omega,\Gamma),\Xi^{s_{1}-2q,\varphi_{1}}_{p}(\Omega,\Gamma)]_{\theta}=\Xi^{s-2q,\varphi}_{p}(\Omega,\Gamma).$
@@
Let us study local properties (up to the boundary $\Gamma$) of the generalized solutions to the elliptic problem. To this end, we introduce local versions of distribution spaces appearing in (3.4). Let $U$ be an open set in $\mathbb{R}^{n}$ such that $\Omega_{0}:=\Omega\cap U\neq\emptyset$, and put $\Gamma_{0}:=\Gamma\cap U$ (the $\Gamma_{0}=\emptyset$ case is admissible). Given $\sigma\in\mathbb{R}$, $\varphi\in\Upsilon$, and $p\in(1,\infty)$, we let $H^{\sigma,\varphi}_{p,\mathrm{loc}}(\Omega_{0},\Gamma_{0})$ denote the linear space of all distributions $u\in\mathcal{D}^{\prime}(\Omega)$ such that $\chi u\in H^{\sigma,\varphi}_{p}(\Omega)$ whenever $\chi\in C^{\infty}(\overline{\Omega})$ and $\mathrm{supp}\,\chi\subset\Omega_{0}\cup\Gamma_{0}$. Analogously, $B^{\sigma,\varphi}_{p,\mathrm{loc}}(\Gamma_{0})$ denotes the linear space of all distributions $v\in\mathcal{D}^{\prime}(\Gamma)$ such that $\chi v\in B^{\sigma,\varphi}_{p}(\Gamma)$ whenever $\chi\in C^{\infty}(\Gamma)$ and $\mathrm{supp}\,\chi\subset\Gamma_{0}$. The operator of the multiplication by any function $\chi\in C^{\infty}(\overline{\Omega})$ (resp., $\chi\in C^{\infty}(\Gamma)$) is bounded on $H^{\sigma,\varphi}_{p}(\Omega)$ (resp., $B^{\sigma,\varphi}_{p}(\Gamma)$), which is known in the $\varphi(\cdot)\equiv 1$ case and then extends by interpolation (Theorems 2.3, 2.7 and Propositions 2.1, 2.8) to the $\varphi\in\Upsilon$ case. Hence, $H^{\sigma,\varphi}_{p}(\Omega)\subset H^{\sigma,\varphi}_{p,\mathrm{loc}}(\Omega_{0},\Gamma_{0})$ and $B^{\sigma,\varphi}_{p}(\Gamma)\subset B^{\sigma,\varphi}_{p,\mathrm{loc}}(\Gamma_{0})$ If $\Omega_{0}=\Omega$ and $\Gamma_{0}=\Gamma$, then these inclusions become equalities.
@@
due to (3.5); i.e., $(f,g)=(A,B)v$ for certain $v\in H^{s,\varphi}_{p}(\Omega)$. Hence, the distribution $u-v\in H^{s-\varepsilon}_{p}(\Omega)$ satisfies (A,B)(uv)=0(A,B)(u-v)=0, which yields $w:=u-v\in C^{\infty}(\overline{\Omega})$ by Theorem 3.1. Thus, $u=v+w\in H^{s,\varphi}_{p}(\Omega)$. We have proved Theorem 3.3 in the case where $\Omega_{0}=\Omega$ and $\Gamma_{0}=\Gamma$, which will help to treat the general case.
@@
Suppose that premise of the application is true. We consider an arbitrary function $\chi\in\Psi$ and choose a function $\eta\in\Psi$ so that $\eta=1$ in the neighbourhood of $\mathrm{supp}\,\chi$. By hypothesis (3.13), we have the inclusion $\chi F\in\Xi^{s-2q,\varphi}_{p}(\Omega,\Gamma)$ where $F:=(A,B)u$. Permuting the differential operator $(A,B)$ and the operator of the multiplication by $\chi$, we obtain
@@
Here, $(A^{\prime},B^{\prime})$ is a differential operator of the form $(A,B)$ with infinitely smooth coefficients such that the orders of all its components are smaller (at least by $1$) than the orders of the corresponding components of the operator $(A,B)$. (Of course, if $\mathrm{ord}\,B_{j}=0$ [i.e., $B_{j}$ is an operator of the multiplication by a function of class $C^{\infty}(\Gamma)$], then the corresponding operator $B_{j}^{\prime}=0$ and we may assume that $\mathrm{ord}\,B_{j}^{\prime}=-1$.)
@@
whatever $p>1$, s>m1+1/ps>m-1+1/p, and $\varphi\in\Upsilon$. This is deduced from the $\varphi(\cdot)\equiv 1$ case by interpolation in an analogous way to that used for (2.26) in the proof of Theorem 2.10, with the condition s>m1+1/ps>m-1+1/p being stipulated by the fact that the orders of all components of $B^{\prime}$ are less than or equal to m1m-1.
@@
We supplement the last theorem with the corresponding a priori estimate of the solution $u$.
@@
If $\Omega_{0}=\Omega$ and $\Gamma_{0}=\Gamma$, then we may take $\chi(t)\equiv\eta(t)\equiv 1$ in (3.20), which means the global a priori estimate of $u$. In this case, estimate (3.20) follows by Peetre’s lemma [53, Lemma 3] from Theorem 3.1 (from the fact that operator (3.4) has finite-dimensional kernel and closed range) and the compact embedding $H^{s,\varphi}_{p}(\Omega)\hookrightarrow H^{s-1,\varphi}_{p}(\Omega)$ (in fact, the continuity of this embedding is enough). In the general case, this estimate is deduced from its global version with the help of (3.15) where $F=(A,B)u$. Namely:
@@
for a certain function parameter $\varphi\in\Upsilon$ subject to (2.32). Then $u\in C^{\ell}(\Omega_{0}\cup\Gamma_{0})$.